# Grammar-based Game Description Generation using Large Language Models

### v1

@@ V-D Quantitative Results
TABLE I: Results of game description generation. Best results shown in bold.
Method Grammar RD GDD Grammar Parsability System Parsability Playability ROUGE
GDG 0.23 0.21 0.19 63.7
GGDG $\checkmark$ 0.32 0.26 0.18 60.8
GGDG + RD $\checkmark$ $\checkmark$ 0.50 0.40 0.22 60.2
GGDG + RD + GDD $\checkmark$ $\checkmark$ $\checkmark$ 0.68 0.40 0.25 60.1
(Oracle Grammar)
GGDG $\checkmark$ 0.40 0.28 0.18 77.1
GGDG + GDD $\checkmark$ $\checkmark$ 0.62 0.39 0.28 78.0
We compare the following methods:
  • Game Description Generation (GDG): Predicts the game description $y$ directly from the demonstration examples and the query $x$ without using the predicted grammar $\hat{G}[y]$.
  • Grammar-based Game Description Generation (GGDG): Generates the grammar $\hat{G}[y]$ required to construct the game description $y$, and then generates $\hat{y}$ based on $\hat{G}[y]$.
  • GGDG + Rule Decoding (RD): Uses the proposed rule decoding in addition to GGDG.
  • GGDG + RD + Game Description Decoding (GDD): Uses both the proposed rule decoding and game description decoding in addition to GGDG.
In addition to these, for GGDG and GGDG + GDD, we also compare the methods that use the oracle grammar $G[y]$ obtained from the ground truth with a parser based on grammar $G$ instead of the generated $\hat{G}[y]$.
The results are summarized in Tab. I. In all metrics, GGDG + RD + GDD shows the highest performance, indicating that introducing rule decoding significantly improves both System Parsability and Playability. Additionally, incorporating game description decoding further enhances Grammar Parsability and Playability. This highlights how both of our decoding methods play an important role in increasing the quality of the generated game descriptions. GDG demonstrates high ROUGE scores but has low scores for Grammar Parsability, System Parsability, and Playability. This suggests that while LLM alone can generate descriptions superficially similar to the ground truth, it is challenging to generate grammatically correct descriptions without accessing Ludii’s grammar. In the case of the oracle grammar, using game description decoding significantly improves Grammar Parsability, System Parsability, and Playability scores. Additionally, compared to not using the oracle grammar, Playability and ROUGE scores are improved, indicating that accurate grammar prediction greatly contributes to the improvement of game descriptions.
V-E Qualitative Results
@@
Figure 8: Qualitative results of game description generation for Tic-Tac-Toe. The columns from left to right are the ground truth, Game Description Generation (GPG), Grammar-based Game Description Generation (GGDG), GGDG with Rule Decoding (GGDG + RD), and GGDG‘+ RD with Game Description Decoding (GGDG + RD + GGD). The middle row shows the game description, and the bottom row shows the results of grammar generation for only the methods that performed grammar generation. Since the grammar results are highly redundant, only the important parts are shown. The highlighted boxes indicate the unnecessary generated parts compared to the ground truth. The further to the right, the fewer such unnecessary parts.
We present the generation results of Tic-Tac-Toe in Fig. 8. The results demonstrate the effectiveness of our proposed framework, particularly the iterative decoding methods. The first row of Fig. 8 displays the results of the game description generation for each method. The far-left column is the ground truth, and the results from the second column onwards are the methods with grammar prediction, rule decoding, and game description decoding added respectively. The second row shows the results of the grammar generation, including only the methods that performed this task. We use Tic-Tac-Die, Order and Chaos, and Unfair as demonstration examples for this test case.
We observe that all results show “Disc” and “Cross” as “X” and “O”. We believe that the LLM is influenced by the “X” and “O” in the query for Tic-Tac-Toe. Tic-Tac-Toe is described in the query $x$ as follows: “Play occurs on a 3x3 grid. One player places an X, the other places an O and players take turns placing their marks in the grid, attempting to get three in a row of their colour.
The generated description of GDG is not Grammar Parsable (nor System Parsable or Playable), and they include the term “count Pips” (highlighted in red) not present in the ground truth. We think that the term “pips” appears because the LLM is influenced by the dice used in the demonstration example of Tic-Tac-Die. The description of Tic-Tac-Die is as follows: “Tic-Tac-Die is played similarly to Tic-Tac-Toe except that players roll a D9 dice each turn to dictate where they move (dice pips show the cell index to move to).
In the game description results of GGDG, we observe that the influence of demonstration examples such as “Pips” has decreased. However, compared to the ground truth, there are extraneous clauses about “move” (highlighted in blue). This occurs because the grammar of “move” predicts a “PASS” clause that is not included in the ground truth.
@@
The grammar results of GGDG + RD show that adding rule decoding aligns them with the grammar of the ground truth. Additionally, the generated game description is System Parsable (but not Playable). However, this description includes the term “do,” which is not in the predicted grammar. This term, found in the demonstration example of Tic-Tac-Die, suggests that the LLM has imitated it.
In the game description results of GGDG + RD + GDD, adding game description decoding reduces the occurrence of expressions not defined in the grammar, such as “do.” Although this description does not fully match the ground truth, it represents the same meaning and is Playable.
VI Limitation
@@
Our proposed framework may have issues with inference time and cost because it repeats LLM inference multiple times. This issue can be mitigated by caching LLM responses, setting appropriate limits on the number of LLM inferences, and improving the prompts.
The evaluation in this paper relies solely on automatically calculable metrics and does not include human evaluations of user experience. The grammar introduced in our framework shows its greatest strength in lower-layer aspects such as Parsability and Playability, rather than in user experience. However, we believe that insights gained from playing the generated games could also lead to further improvements.
VII Broader Impact
@@
Our research is related to copyright issues concerning input and output data. It is common practice for LLMs to be trained using copyrighted data [11]. Since our framework does not include the training process, it does not directly cause this issue. However, when using our framework, attention should be paid to the training data of LLMs. Our framework automatically generates game descriptions without human intervention. The output generated by our framework may not be copyrighted. Developing our framework into an interactive approach with human designers may allow the final output to qualify for copyright.
The generation process of LLMs is opaque, and in our research, the process for each LLM call is similarly unclear. However, our grammar-based iterative decoding method breaks down the generation process into multiple steps, which should enhance our understanding of it.
@@
VIII Conclusions
In this paper, we proposed an automated game design framework using Large Language Models (LLMs) with the Ludii grammar as our Game Design Language grammar. Our framework generates Ludii game descriptions based on in-context learning with LLMs. The generation process is performed in two stages: generating the grammar necessary for the game description, and generating the game description itself. Our proposed approach enables the LLMs to capture the features of the game description generation task with fewer demonstration examples. We further propose iterative decoding methods specialized for both the grammar and the game descriptions, which improve the grammatical correctness of the generated game descriptions. Extensive experimental results show the effectiveness of our framework in the game description generation task.

### v2

@@ VI-C Ablation Study
TABLE IV: Comparison of the iteration limits for Rule Decoding. Game Description Decoding is not used (i.e., GGDG w/o GDD). The best results are in bold.
[…]
10 50.0$\pm$1.0 44.7$\pm$2.0 61.5$\pm$0.1 0.58$\pm$0.02
20 51.3$\pm$3.5 46.7$\pm$2.4 61.3$\pm$0.2 0.55$\pm$0.02
30 47.7$\pm$2.7 44.0$\pm$3.5 62.1$\pm$0.2 0.58$\pm$0.03
TABLE V: Comparison of the iteration limits for Game Description Decoding. The number of iterations for Rule Decoding is fixed at 20. The best results are in bold.
[…]
5 60.0$\pm$3.2 53.7$\pm$2.8 61.1$\pm$0.2 0.50$\pm$0.02
10 64.0$\pm$1.5 56.7$\pm$2.3 60.5$\pm$0.6 0.46$\pm$0.03
20 64.7$\pm$0.3 57.3$\pm$1.9 60.6$\pm$0.2 0.47$\pm$0.01
Iteration Limit for Decoding. We investigate the impact of iteration limits in our decoding approach. First, Tab. IV shows the results of rule decoding. When the iteration limit is set to 30, we find that the Compilability and Functionality scores are the lowest. This may be because increasing the iteration limit leads to the inclusion of unnecessary rules, potentially degrading performance. When the iteration limit is set to 20, Compilability, Functionality, and NCD scores show the best results. Therefore, we set the iteration limit to 20 for subsequent experiments.
Tab. V shows the results of game description decoding. When the iteration limit is set to 5, we find that the Compilability and Functionality scores are the lowest. This is likely due to insufficient iterations for improving the game description. When the iteration limit is set to 10 or 20, we observe minimal differences across most metrics. When the iteration limit is 10, NCD is lowest at 0.46 and the computational cost is also low, therefore we set the iteration limit to 10 in other experiments.
TABLE VI: Comparison of game categories in demonstration examples. “Same” category indicates that examples are from the same category as the test instance, while “Cross” category indicates that they are from different categories. The best results are in bold.
[…]
Cross Category
GDG 5.3$\pm$0.3 3.7$\pm$0.7 46.0$\pm$0.0 0.97$\pm$0.00
GGDG 36.0$\pm$0.6 24.6$\pm$0.3 42.3$\pm$0.1 0.79$\pm$0.00
Same Category
GDG 27.0$\pm$1.2 26.3$\pm$0.7 63.5$\pm$0.6 0.75$\pm$0.01
GGDG 64.0$\pm$1.5 56.7$\pm$2.3 60.5$\pm$0.6 0.46$\pm$0.03
Game Category of Demonstration Examples. We investigate how the category of demonstration examples affects game performance. In our proposed method, we use instances from the same category as the test instances for demonstration examples (Same Category). We compare this with using instances from different categories (Cross Category), and summarize the results in Tab. VI. In Cross Category, demonstration examples are randomly selected from all games except for the category of the test instance. We find that obtaining demonstration examples from the same category as test instances significantly improves performance across all evaluation metrics.
VI-D Impact of Game Characteristics and Model Configurations
TABLE VII: Comparison of test game lengths. 300 - 500 indicates that the evaluation is performed on games with game description token lengths between 300 and 500.
[…]
0 - 300
GDG 27.0$\pm$1.2 26.3$\pm$0.7 63.5$\pm$0.6 0.75$\pm$0.01
GGDG 64.0$\pm$1.5 56.7$\pm$2.3 60.5$\pm$0.6 0.46$\pm$0.03
300 - 500
GDG 41.0$\pm$2.1 39.0$\pm$1.0 59.7$\pm$0.2 0.68$\pm$0.01
GGDG 54.3$\pm$2.2 48.7$\pm$1.3 59.4$\pm$0.5 0.59$\pm$0.01
500 - 1000
GDG 36.0$\pm$2.3 32.7$\pm$2.6 49.9$\pm$0.3 0.74$\pm$0.02
GGDG 33.3$\pm$1.2 29.7$\pm$0.9 47.3$\pm$0.6 0.77$\pm$0.00
Game Description Length. We investigated the impact of game description length on performance. We compared three groups based on token length: 0-300, 300-500, and 500-1,000. Token lengths were calculated using the Llama-3-8B-Instruct tokenizer. The results are summarized in Tab. VII. For 0-300 and 300-500 tokens, GGDG outperforms GDG in Compilability, Functionality, and NCD metrics, demonstrating its ability to generate more grammatically accurate game descriptions. Meanwhile, GGDG showed declining performance across all metrics as token length increased, dropping from 64.0 to 33.3, Functionality from 56.7 to 29.7, ROUGE from 60.5 to 47.3, and NCD worsening from 0.46 to 0.77. For 500-1,000 tokens, GDG slightly outperformed GGDG across all metrics. These results suggest that GGDG struggles with long game descriptions. We discuss this limitation in Sec VII.
TABLE VIII: Comparison of test instance categories.
[…]
board/race
GDG 10.0$\pm$5.8 10.0$\pm$5.8 58.2$\pm$0.2 0.90$\pm$0.06
GGDG 60.0$\pm$5.8 60.0$\pm$5.8 58.8$\pm$0.2 0.42$\pm$0.05
board/sow
GDG 38.9$\pm$5.6 38.9$\pm$5.6 78.8$\pm$1.0 0.75$\pm$0.06
GGDG 55.6$\pm$5.6 55.6$\pm$5.6 81.3$\pm$0.9 0.51$\pm$0.06
puzzle
GDG 7.4$\pm$1.9 7.4$\pm$1.9 51.3$\pm$0.5 0.98$\pm$0.02
GGDG 61.1$\pm$6.4 59.3$\pm$6.7 52.1$\pm$0.7 0.57$\pm$0.12
board/space/line
GDG 26.9$\pm$0.9 24.4$\pm$0.5 64.2$\pm$0.3 0.78$\pm$0.00
GGDG 68.7$\pm$0.9 63.7$\pm$1.3 62.4$\pm$0.4 0.42$\pm$0.02
board/war
GDG 57.7$\pm$4.4 57.7$\pm$4.4 71.0$\pm$0.1 0.45$\pm$0.04
GGDG 71.8$\pm$7.1 62.8$\pm$9.0 71.8$\pm$0.3 0.41$\pm$0.09
Game Category of Test Games. We investigate the impact of game categories of test instances. We compare GDG and GGDG across five categories: racing games (board/race), mancala games (board/sow), puzzle games (puzzle), line games (board/space/line), and war games, including capture games (board/war). The demonstration examples are from the same category as the test instances. We use instances with game description token lengths of 300 or less as test games.
The results are summarized in Tab. VIII. Across all categories, our proposed GGDG outperforms GDG in terms of Compilability, Functionality, and NCD metrics, which is consistent with our previous experimental results. For ROUGE scores, while the superior method varies by category, the difference between GDG and GGDG remains within 2 points across most categories.
In comparisons across categories, GDG’s performance shows significant variation, with the puzzle category showing the lowest scores (Compilability/Functionality: 7.4, NCD: 0.98) for GDG and the largest improvement margin (+53.7/+51.9 and -0.41 respectively) when using GGDG. Conversely, the board/war categories showed GDG’s highest performance levels (Compilability/Functionality: 57.7, NCD: 0.45) and the smallest margin of improvement with GGDG (+14.1/+5.1 and -0.04 respectively). This variation may be attributed to games in the puzzle category having more complex logic compared to other categories, making it difficult for LLMs to learn grammatically correct patterns from demonstration examples, or possibly due to limited knowledge about puzzle category games obtained during pre-training. Meanwhile, GGDG showed smaller performance variations between categories compared to GDG, with scores ranging from 55.6 to 71.8 for Compilability, 55.6 to 63.7 for Functionality, and 0.41 to 0.57 for NCD, demonstrating more stable performance.
TABLE IX: Comparison of LLM models. The best results are in bold.
[…]
Llama-3.2-3B-Instruct
GDG 18.7$\pm$0.7 17.3$\pm$0.3 59.2$\pm$0.3 0.83$\pm$0.00
GGDG 28.0$\pm$1.5 25.7$\pm$2.0 55.4$\pm$0.1 0.76$\pm$0.02
Llama-3-8B-Instruct
GDG 27.0$\pm$1.2 26.3$\pm$0.7 63.5$\pm$0.6 0.75$\pm$0.01
GGDG 64.0$\pm$1.5 56.7$\pm$2.3 60.5$\pm$0.6 0.46$\pm$0.03
gpt-4o
GDG 27.7$\pm$0.3 27.7$\pm$0.3 67.7$\pm$0.1 0.74$\pm$0.00
GGDG 70.7$\pm$0.9 59.3$\pm$0.9 68.2$\pm$0.0 0.44$\pm$0.01
LLM Models. We compare the performance of different LLM models. We primarily used Llama-3-8B-Instruct in our experiments and compare its performance to that of Llama-3.2-3B-Instruct, a smaller LLM model, and gpt-4o, a more powerful LLM model. The results are summarized in Tab. IX. GGDG improves Compilability, Functionality, and NCD scores over GDG across all models. With gpt-4o, GGDG also improves the ROUGE score. Comparing across models, the 8B model shows significant improvements in all evaluation metrics compared to the 3B model. For example, with GGDG, upgrading from the 3B model to the 8B model improves Compilability by +36.0, Functionality by +31.0, ROUGE by +5.1, and NCD by -0.30. gpt-4o achieves the highest scores across all evaluation metrics. The improvements over the 8B model are +6.7 in Compilability, +2.6 in Functionality, +7.7 in ROUGE, and -0.02 in NCD. We observe that the improvements in Compilability, Functionality, and NCD are larger when moving from the 3B model to the 8B model compared to moving from the 8B model to gpt-4o. This suggests that while increasing model size brings significant performance improvements, the room for improvement becomes limited between already high-performing models. Therefore, if higher performance is desired, selecting powerful models like gpt-4o is effective. In contrast, if there are resource constraints, such as API costs or computational resource limitations, the 8B model can be considered a balanced choice between performance and cost.
VI-E Qualitative Results
Figure 8: Comparison of generation results with baseline methods for Latin Square. Since the board size is not specified in the query, it varies depending on the generated results. The ground truth is set to size 5 as an example.
Comparison with Baseline Methods. We conduct a qualitative analysis comparing SFT+GDG, which is the most promising among the baseline methods, with our proposed methods, GGDG and SFT+GGDG. Figure 8 shows the generation results for Latin Square. Latin Square is a puzzle where players place numbers in an n×$\times$×n grid such that the same number does not repeat in each row and column.
The result of SFT+GDG is neither compilable nor functional due to the part shown in the red frame: “(forAll Region (== 0 (state at:(to)))”. This issue stems mainly from two grammatical problems. First, while the Ludii grammar defines that a “forAll” clause must be followed by a “puzzleElementType” (either “Cell”, “Edge”, “Vertex”, or “Hint”), an undefined terminal symbol, “Region”, is used. Second, the expression “(== 0 (state at:(to)))” cannot be parsed because the operator “==” is not defined in the Ludii grammar.
In contrast, GGDG’s result is both compilable and functional, generating a game description almost identical to the ground truth. In particular, it shows significant improvement in that the result contains no grammatical errors. However, in the blue frame, “except:0” (a constraint requiring all values to be different except for index 0) is added after “all Different”, and this rule is not included in the game rules of Latin Square.
Furthermore, the SFT+GGDG result shows more improvement and is consistent with the Latin Square game rules. The “is Full” rule, which requires all cells to be filled, is redundant as it is already included in the content of the “all Different” rule, but it does not contradict the Latin Square rules. These results suggest that combining SFT and GGDG improves performance.
@@ VI-E Qualitative Results
Figure 9: Impact of GGDG components on generation results for Tic-Tac-Toe. Components are added from left to right, with the ground truth on the far right. The middle row shows the game description, and the bottom row shows the results of grammar generation for only the methods that performed grammar generation. Since the grammar results are highly redundant, only the important parts are shown.
GGDG Components. We investigate the effects of GGDG components through qualitative analysis. Figure 9 shows the generation results for Tic-Tac-Toe. The GDG results are not compilable due to the “count Pips” shown in the red frame. The appearance of the term “Pips” occurs because the LLM is influenced by the dice in the Tic-Tac-Die game from the demonstration examples. Tic-Tac-Die is explained as follows: “Tic-Tac-Die is played similarly to Tic-Tac-Toe except that players roll a D9 dice each turn to dictate where they move (dice pips show the cell index to move to).”
While the pips disappear in the results of GGDG w/o RD, GDD, they are still not compilable. This is due to the extra clause about “PASS” in the blue frame. This occurs because the predicted “move” grammar contains a “PASS” clause not included in the ground truth.
@@
The grammar results of GGDG w/o GDD show that by adding rule decoding, they match the ground truth grammar. The generated game description is functional. However, this description includes “do,” which is not in the predicted grammar. The presence of this term, which appears in the Tic-Tac-Die demonstration example, suggests that the LLM is mimicking it.
In the GGDG results, adding game description decoding reduces the occurrence of expressions not defined in the grammar, such as “do”. This game description is functional and closest to the ground truth among the compared methods. However, since the condition “(is line 3)” is applied equally to player1 and player2 as shown in the purple frame, it cannot accurately determine which player has won. This suggests there is still room for improvement in GGDG.
Figure 10: Comparison of generation results for Bara Guti (Bihar) as a failure example. Some parts of the game description are omitted due to space constraints. The board appears in the upper right of each game description. Due to space constraints, GGDG’s results display a wedge of 4 instead of 8.
Analysis of Failure Cases. We examine the analysis results of Bara Guti (Bihar) as an example of a failure case, as shown in Fig. 10. This game uses a board that has three concentric circles, with four diameters dividing it into eight equal sections, with counters placed in each section. In analyzing the generated results, we note that ”Markers” and ”Counters” refer to the same type of piece.
While SFT+GDG generated compilable output, it did not produce functional results due to the parts highlighted in red in Fig. 10. “concentric 1 8 8” specifies the number of cells in each concentric ring sequentially, defining a total of 17 cells. However, this description becomes non-functional because lines 19 and 21 specify marker initial positions with indices of 17 or greater. Additionally, the specified initial positions differ from the ground truth.
In contrast, while GGDG’s generated results demonstrate functionality, the definitions of the board and counter initial positions differ from the ground truth. Specifically, it generates a board by connecting two triangles on lines 19 and 22, and places counters at positions expanded two steps from the bottom using the expression “expand (sites Bottom) steps:2” (highlighted in blue). This difference is presumed to be influenced by the existing demonstration example of Lau Kata Kati: “(merge (wedge 4) (shift 0 3 (rotate 180 (wedge 4))))”.
Finally, while SFT+GGDG’s results maintain functionality and return to the concentric circle board structure, the number of concentric circles, cell count, and Marker initial positions are incorrectly predicted, highlighted in green. These limitations can be attributed to LLM’s lack of ability to understand spatial structures like boards [72]. We discuss these limitations of spatial understanding in Sec VII.
VII Limitations and Discussion
As shown in Tab. VII, GGDG’s performance decreases with longer game descriptions. It is known that LLM performance degrades with longer input sequences [73]. We believe that GGDG struggles with processing longer game descriptions because its input token length is increased by the grammar, compared to GDG. The development of LLMs capable of handling longer contexts is ongoing [74, 75], and newer LLM models may help mitigate this issue.
As shown in Fig. 10, it is difficult to compensate for LLMs’ lack of spatial understanding capabilities through SFT. It has been found that LLMs with 70B parameters demonstrate superior capabilities compared to models with 7B or 13B parameters in this regard [72]. Using a larger LLM, such as the 70B parameter models, is a possible option to increase such capabilities. However, it should be noted that SFT for large-parameter LLMs requires substantial computational resources, which can be problematic in other ways.
@@ VII Limitations and Discussion
Our proposed framework may have issues with inference time and cost because it repeats LLM inference multiple times. Specifically, the inference time per game averages 7.7 seconds for GDG and 143.9 seconds for GGDG. This issue can be mitigated by caching LLM responses, setting appropriate limits on the number of LLM inferences, and improving the prompts. Additionally, given recent advances in inference acceleration techniques [76], these constraints are expected to be lessened as newer techniques are developed.
The evaluation in this paper relies solely on automatically calculable metrics and does not include human evaluations. We believe that human evaluation could capture errors that automated evaluation cannot fully cover, potentially leading to further improvements and leads for future research directions.
VIII Broader Impact
@@
Our research is related to copyright issues concerning input and output data. It is common practice for LLMs to be trained using copyrighted data [39]. Since our framework does not include the training process, it does not directly cause this issue. However, users should consider the training data of LLMs when employing our framework. Our approach automatically generates game descriptions without human intervention, and these outputs may not be copyrighted. Developing our framework into an interactive approach with human designers may allow the final output to qualify for copyright.
The generation process of LLMs is opaque, and there are studies examining whether LLMs’ black-box nature hinders the transparency of PCG [44]. In our research as well, the process of each LLM call remains similarly unclear. However, our grammar-based iterative decoding method breaks down the generation process into multiple steps, which should enhance our understanding of it.
@@
IX Conclusions
We have proposed a novel framework that combines large language models with game description languages for generating game descriptions from text. Our approach integrates GDL grammar into the generation process, enabling the creation of structurally coherent game descriptions. By introducing iterative refinement decoding methods specialized for both grammar generation and game description generation, we have seen improvements in the grammatical accuracy of game descriptions. Extensive experimental results demonstrate that our framework is effective in improving grammatical accuracy within game description generation. Future research may explore fine-tuning specialized LLMs for each subtask of rule and game description completion performed at each step of iterative refinement decoding, as well as the utilization of more efficient and powerful LLMs. Additionally, it would be valuable to qualitatively validate the findings of our approach beyond the Ludii dataset through conducting user studies with actual humans.
Acknowledgments
This work was supported by JST, ACT-X Grant Number JPMJAX23CE, Japan.